A system of equations is given below, 𝑑π‘₯ + 2𝑦 + 3𝑧 = π‘Ž 2π‘₯ + 3𝑦 βˆ’ 𝑑𝑧 = 𝑏 3π‘₯ + 5𝑦 + (𝑑 + 1)𝑧 = 𝑐 Where 𝑑 is an integer and π‘Ž, 𝑏, 𝑐 are real constants. The system does not have a unique solution, but it is consistent. Show that π‘Ž + 𝑏 = 𝑐.


1
Expert's answer
2021-09-28T10:41:32-0400

System of equations are: 𝑑π‘₯ + 2𝑦 + 3𝑧 = π‘Ž, 2π‘₯ + 3𝑦 βˆ’ 𝑑𝑧 = 𝑏, 3π‘₯ + 5𝑦 + (𝑑 + 1)𝑧 = 𝑐.

Now, "\begin{vmatrix}
 t & 2 & 3 \\
 2 & 3 &-t \\
3 & 5 & (t+1)
\end{vmatrix}=0" [Β System does not have a unique solution, but it is consistent ]

"t[3t+3+5t]-2[2t+2+3t]+3[10-9]=0\\
\Rightarrow 8t^2+3t-10t-4+3=0\\
\Rightarrow 8t^2-7t-1=0\\
\Rightarrow 8t^2-8t+t-1=0\\
\Rightarrow (8t+1)(t-1)=0\\
\therefore t=1,\frac{-1}{8}"

But ""t"" is an integer.

So, "t=1"

Now, system of equations are: π‘₯ + 2𝑦 + 3𝑧 = π‘Ž, 2π‘₯ + 3𝑦 βˆ’ 𝑧 = 𝑏, 3π‘₯ + 5𝑦 + 2𝑧 = 𝑐.

So, adding first 2 equations, we get: 3π‘₯ + 5𝑦 + 2𝑧 = π‘Ž+𝑏


Comparing this equation with third equation, we get:

"a+b=c"

Hence Proved.